OpenStudy (secret-ninja):

Help please. Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y = -6x - 15 y = -6x - 20 y = -6x + 15 y = -6x + 20

OpenStudy (amistre64):

the equation will prove true for the given point, since they all of -6x in common.

OpenStudy (amistre64):

if we generalize it: y = m(x-a) + b y = mx -ma + b so the constant at the end is simply equal to -ma+b

OpenStudy (secret-ninja):

I just have to find out what "b" is, and if it is positive or negative right?

OpenStudy (amistre64):

pretty much yeah. if you dont recall the point slope format, then the simplest process is to insert the points x and y values to see what is true

OpenStudy (secret-ninja):

Okay i'm going to try and get the answer Gimme a minute.

OpenStudy (amistre64):

for example y = -6x - 20, (-2,-3) -3 = -6(-2) - 20 -3 = 18-20 -3 = -2 is not good

OpenStudy (amistre64):

i dont see a B option, so im not sure which one you are referinng to.

OpenStudy (secret-ninja):

Crap sorry I meant A. y = -6x - 15 I didn't solve them in the same order as listed.

OpenStudy (secret-ninja):

I think it is A. lol

OpenStudy (amistre64):

A is fine :) and to correct my inability to multiply lol y = -6x - 20, (-2,-3) -3 = -6(-2) - 20 -3 = 12-20 -3 = -8 is not good

OpenStudy (secret-ninja):

Okay thank you for your help! :)